# Mirror Descent Algorithms with Nearly Dimension-Independent Rates for Differentially-Private Stochastic Saddle-Point Problems

### v1

@@ 2 Preliminaries
Given $x\in\mathbb{R}^{d}$, we denote the $1$-norm and $\infty$-norm by $\|x\|_{1}=\sum_{j\in[d]}|x_{j}|$ and $\|x\|_{\infty}=\max_{j\in[d]}|x_{j}|$, respectively. A differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $L_{0}$-Lipschitz w.r.t. $\|\cdot\|_{1}$ over a set $\mathcal{X}$ if $\|\nabla F(x)\|_{\infty}\leq L_{0}$ for all $x\in\mathcal{X}$. It is $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$ if $\|\nabla F(x^{1})-\nabla F(x^{2})\|_{\infty}\leq L_{1}\|x^{1}-x^{2}\|_{1}$ for all $x^{1},x^{2}\in\mathcal{X}$. We say a function is first-order-smooth if it is $L_{0}$-Lipschitz and $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$. We say a function is second-order-smooth if it is first-order-smooth and its partial derivates are $L_{2}$-smooth w.r.t $\|\cdot\|_{1}$. [^5] Note that this assumption is milder than the typical second-order smoothness assumption in optimization, and it is satisfied by the widely used quadratic and logistic losses. For $a,b:\mathbb{R}^{p}\mapsto\mathbb{R}$, we say $a\lesssim b$ if there exists an absolute constant $C>0$ such that for all $x\in\mathbb{R}^{p}$, $a(x)\leq Cb(x)$. When $a\lesssim b$ and $b\lesssim a$, we write $a\eqsim b$. We denote by $e_{i}$ the $i$-th canonical vector (the dimension being clear from context). The standard simplex in $\mathbb{R}^{d}$ is the set $\Delta_{d}=\{x\in\mathbb{R}^{d}:\|x\|_{1}=1,x_{j}\geq 0\text{ for all }j\in[d]\}$. A vector $x\in\Delta$ induces a probability distribution over the canonical vectors, which we denote by $P_{x}$. Hence, $\hat{x}\sim P_{x}$ means that $\hat{x}$ takes the value $e_{i}$ with probability $x_{i}$. Note that if $\hat{x}\sim P_{x}$, then $\mathbb{E}[\hat{x}]=x$. Finally, we define the notion of differential privacy.
Definition 2.1 (Differential Privacy Dwork et al. (2006a)).
A randomized algorithm $\mathcal{A}:\mathcal{Z}^{n}\mapsto\mathcal{X}$ is $(\varepsilon,\delta)$-differentially private (abbreviated $(\varepsilon,\delta)$-DP) if for any pair of datasets $S$ and $S^{\prime}$ differing in at most one data point (denoted by $S\sim S^{\prime}$) and any event $\mathcal{E}\subseteq\mathcal{X}$, $\mathbb{P}[\mathcal{A}(S)\in\mathcal{E}]\leq e^{\varepsilon}\mathbb{P}[\mathcal{A}(S^{\prime})\in\mathcal{E}]+\delta$. When $\delta=0$, we denote the $(\varepsilon,0)$-DP property as $\varepsilon$-DP.
@@ 3 Maurey Sparsification with Functional Value Approximation
The classical Maurey Sparsification Lemma (Pisier, 1980-1981) is an application of the probabilistic method that leads to an approximate and (nearly) dimension-independent version of Carathéodory’s Theorem (Carathéodory, 1911). While this method is well-known to apply with smooth norms, it appears that for norms such as $\|\cdot\|_{1}$ such approximations are out of reach. [^6] In fact, $\|\cdot\|_{1}$ is as far as being smooth as possible, in the sense of having trivial Rademacher type (Pisier, 2016). Furthermore, polynomial in $d$ mean estimation lower bounds with respect to $\|\cdot\|_{1}$ are known (Waggoner, 2015). We circumvent this limitation by directly working with function value approximations. This introduces its own additional challenges which we address in Section 4. Proofs in this section can be found in Appendix D.
We start with a simple adaptation of Maurey’s method for function approximation in the general smooth case. A similar proof appears in (Acharya et al., 2019).

### v2

@@ 1.3 Preliminaries
Given $x\in\mathbb{R}^{d}$, we denote the $1$-norm and $\infty$-norm by $\|x\|_{1}=\sum_{j\in[d]}|x_{j}|$ and $\|x\|_{\infty}=\max_{j\in[d]}|x_{j}|$, respectively. A differentiable function $F:\mathbb{R}^{d}\to\mathbb{R}$ is $L_{0}$-Lipschitz w.r.t. $\|\cdot\|_{1}$ over a set $\mathcal{X}$ if $\|\nabla F(x)\|_{\infty}\leq L_{0}$ for all $x\in\mathcal{X}$. It is $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$ over $\mathcal{X}$ if $\|\nabla F(x^{1})-\nabla F(x^{2})\|_{\infty}\leq L_{1}\|x^{1}-x^{2}\|_{1}$ for all $x^{1},x^{2}\in\mathcal{X}$. We say a function is first-order-smooth if it is $L_{0}$-Lipschitz and $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$. We say a function is second-order-smooth if it is first-order-smooth and its partial derivatives are $L_{2}$-smooth w.r.t $\|\cdot\|_{1}$. [^4] Note that this assumption is milder than the typical second-order-smoothness assumption in optimization, and it is satisfied by the widely used quadratic and logistic losses. For $a,b:\mathbb{R}^{p}\mapsto\mathbb{R}$, we say $a\lesssim b$ if there exists an absolute constant $C>0$ such that for all $x\in\mathbb{R}^{p}$, $a(x)\leq Cb(x)$. When $a\lesssim b$ and $b\lesssim a$, we write $a\eqsim b$. We denote by $e_{i}$ the $i$-th canonical vector (the dimension being clear from context). The standard simplex in $\mathbb{R}^{d}$ is the set $\Delta_{d}=\{x\in\mathbb{R}^{d}:\|x\|_{1}=1,x_{j}\geq 0\text{ for all }j\in[d]\}$. A vector $x\in\Delta_{d}$ induces a probability distribution over the canonical vectors, which we denote by $P_{x}$. Hence, $\hat{x}\sim P_{x}$ means that $\hat{x}$ takes the value $e_{i}$ with probability $x_{i}$. Note that if $\hat{x}\sim P_{x}$, then $\mathbb{E}[\hat{x}]=x$. Given $f:{\cal X}\times{\cal Y}\mapsto\mathbb{R}$, we let $\Phi(x,y)=\big(\nabla_{x}f(x,y),-\nabla_{y}f(x,y)\big)$ be saddle operator (a.k.a. the monotone operator associated to the saddle-point problem) [Rockafellar:1970]. We denote $\ell=\log(d_{x})+\log(d_{y})$, where $d_{x},d_{y}$ are the ambient dimensions of the optimization variables in the SSP problem.
1.4 Motivating Examples
As motivating examples for the DP-SSP problem (1), we present two problems of interest in privacy-preserving machine learning and data analysis that fit into our problem setup.
DP Synthetic Data Generation
Consider a finite sample space $\mathcal{Z}=\{z_{1},...,z_{|\mathcal{Z}|}\}$ and an unknown distribution $\mathcal{D}$ supported on $\mathcal{Z}$. Suppose we have a dataset $S=\{z_{i_{1}},...,z_{i_{n}}\}\overset{iid}{\sim}\mathcal{D}$ and a set of queries $\mathcal{Q}=\{q:\mathcal{Z}\to[-1,1]\}$. Defining $q(S)=\frac{1}{|S|}\sum_{z\in S}q(z)$ and $q(\mathcal{D})=\mathbb{E}_{z\sim\mathcal{D}}[q(z)]$, our goal is to construct a synthetic dataset $\tilde{S}$ such that $\tilde{S}$ is private with respect to $S$ and $\max_{q\in{\cal Q}}|q(\tilde{S})-q(\mathcal{D})|$ is as small (in expectation) as possible. Using a linear relaxation for the discrete maximum over $\cal Q$} and viewing each query $q$ as vector $(q(z))_{z\in\mathcal{Z}}$, we obtain a bilinear SSP formulation as follows:
(3) $\min_{x\in\Delta_{|\mathcal{Z}|}}\max_{y\in\Delta_{|\mathcal{Q}|}}\mathbb{E}_{z\sim\mathcal{D}}\Big[\sum_{j\in[|\mathcal{Q}|]}y_{j}(q_{j}(z)-\langle q_{j},x\rangle)\Big].$
While our Algorithm 1, proposed in Section 3.1, can solve this problem, applying it directly would result in a slightly suboptimal dependence on the histogram size, $|{\cal Z}|$. However, a variant of it that leverages the specific structure of this problem attains the optimal rates [lower_bound_DP_query_answering]. More concretely, the following theorem holds.
Theorem 1.1.
Let $0<\delta<1$, $0<\varepsilon<8\log(1/\delta)$. If $|\mathcal{Q}|\leq|\mathcal{Z}|^{C}$, then a variant of Algorithm 1 constructs an $(\varepsilon,\delta)$-DP synthetic dataset $\tilde{S}$ of $n$ samples, such that
$\mathbb{E}\Big[\max_{q\in\mathcal{Q}}|\mathbb{E}_{z\sim\mathcal{D}}[q(z)]-q(\tilde{S})|\Big]\lesssim\sqrt{\frac{\log(|{\cal Q}|)}{n}}+(1+C)\frac{(\log(1/\delta)\log(|\mathcal{Z}|))^{1/4}\sqrt{\log(|{\cal Q}|)}}{\sqrt{n\varepsilon}}.$
Furthermore, this rate is optimal up to constant factors.
Minimizing the maximal loss under DP
Consider a collection of convex losses $f_{i}:\mathcal{X}\times\mathcal{Z}\to[-B,B]$, $i\in[d_{y}]$. Suppose $\mathcal{X}$ is a polytope. We are interested in minimizing $\max_{i\in[d_{y}]}\{F_{i}(x):=\mathbb{E}_{z\sim\mathcal{D}}[f_{i}(x;z)]\}$. This can be formulated as follows:
The maximal loss formulation encompasses applications to fairness [Jagielski:2019, Fair_FL, Lowy:2023] and robustness [Mohajerin:2018] under privacy constraints. As a remark, if the functions $(f_{i})_{i\in[d_{y}]}$ are $L_{0}$-Lipschitz and $L_{1}$-smooth on $x$ w.r.t $\|\cdot\|_{1}$, then the problem objective $f(x,y;z)=\sum_{i\in[d_{y}]}y_{i}f_{i}(x;z)$ is convex-concave, $\max\{L_{0},B\}$-Lipschitz, $\max\{L_{0},L_{1}\}$-smooth on $(x,y)$. If the functions $(f_{i})_{i\in[d_{y}]}$ are also $L_{2}$-second-order-smooth on $x$, then $f(x,y)$ is $\max\{L_{1},L_{2}\}$-second-order-smooth on $(x,y)$. Examples of convex losses that are Lipschitz, smooth and second-order-smooth over compact sets are the quadratic and the logistic loss.
@@ 2 Maurey Sparsification with Functional Value Approximation
The classical Maurey Sparsification Lemma [Pisier:1980] is an application of the probabilistic method that leads to an approximate and (nearly) dimension-independent version of Carathéodory’s Theorem [Carathodory:1911]. While this method is well-known to apply with smooth norms, it appears that for norms such as $\|\cdot\|_{1}$ such approximations are out of reach. [^5] In fact, $\|\cdot\|_{1}$ is as far as being smooth as possible, in the sense of having trivial Rademacher type [Pisier:2016]. Furthermore, polynomial in $d$ mean estimation lower bounds with respect to $\|\cdot\|_{1}$ are known [Waggoner:2015]. We circumvent this limitation by directly working with function value approximations. These results will be crucial for the analysis of our algorithms in the forthcoming sections.
We start with a simple adaptation of Maurey’s method for function approximation in the general smooth case. A similar proof appears in [Acharya:2019]. Throughout this section we denote $\bar{x}^{T}=\sum_{t=1}^{T}\frac{x^{t}}{T}$ and $\bar{a}^{T}=\sum_{t=1}^{T}\frac{a^{t}}{T}$ unless explicitly said otherwise.
@@
Let $(\mathbf{E},\|\cdot\|)$ be a normed space, and let $F$ be a $L_{1}$-smooth function w.r.t. $\|\cdot\|$. Let $x^{1},...,x^{T}$ be a collection of deterministic vectors, $(\lambda_{1},\ldots,\lambda_{T})\in\Delta_{T}$, and $a^{1},...,a^{T}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$. Let $\bar{x}^{T}=\sum_{t=1}^{T}\lambda_{t}x^{t}$ and $\bar{a}^{T}=\sum_{t=1}^{T}\lambda_{t}a^{t}$. Then $\big|\mathbb{E}\left[F\left(\bar{a}^{T}\right)-F\left(\bar{x}^{T}\right)\right]\big|\leq\frac{L_{1}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\mathbb{E}\left[\left\|a^{t}-x^{t}\right\|^{2}\right]$.
Proof 2.2.
Let $\alpha_{t}=\Big\langle\nabla F\Big(\sum_{k=t}^{T}\lambda_{k}x^{k}+\sum_{k=1}^{t-1}\lambda_{k}a^{k}\Big),\lambda_{t}(a^{t}-x^{t})\Big\rangle$. Noting that $F(\bar{a}^{T})=F(\bar{x}^{T}+\bar{a}^{T}-\bar{x}^{T})$, and using repeatedly the $L_{1}$-smoothness of $F$, we have
$F(\bar{a}^{T})$ $\leq F(\bar{x}^{T}+\bar{a}^{T-1}-\bar{x}^{T-1})+\alpha_{T}+\frac{L_{1}\lambda_{T}^{2}\left\|a^{T}-x^{T}\right\|^{2}}{2}$
$\ldots$ $\leq F\left(\bar{x}^{T}\right)+\sum_{t\in[T]}\alpha_{t}+\frac{L_{1}}{2}\sum_{t\in[T]}\lambda_{t}^{2}\|a^{t}-x^{t}\|^{2}.$
Clearly $(\alpha_{t})_{t\in[T]}$ is a martingale difference sequence. Hence, $\mathbb{E}\big[\sum_{t\in[T]}\alpha_{t}\big]=0$ and consequently $\mathbb{E}[F(\bar{a}^{T})-F(\bar{x}^{T})]\leq\frac{L_{1}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\mathbb{E}\left[\left\|a^{t}-x^{t}\right\|^{2}\right]$. Since F-F is also $L_{1}$-smooth, the same idea yields $\mathbb{E}[F(\bar{x}^{T})-F(\bar{a}^{T})]\leq\frac{L_{1}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\mathbb{E}\left[\left\|a^{t}-x^{t}\right\|^{2}\right]$.
We now instantiate Lemma 2.1 for the space $(\mathbb{R}^{d},\|\cdot\|_{1})$ with $\lambda_{i}=\frac{1}{T}$, $i\in[T]$. In later sections we will see that $x^{t}\in\Delta_{d}$ represents the $t$-th iterate of our algorithm and $a^{t}\sim P_{x^{t}}$ a private, unbiased estimator of $x^{t}$.
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be a $L_{1}$-smooth function w.r.t. $\|\cdot\|_{1}$. Let $x^{1},...,x^{T}\in\Delta_{d}$ be deterministic vectors and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for all $t\in[T]$. Then $\left|\mathbb{E}\left[F\left(\bar{a}^{T}\right)-F\left(\bar{x}^{T}\right)\right]\right|\leq\frac{2L_{1}}{T}$.
The above result bounds the bias incurred by evaluating gradients at sampled vertices. More concretely, the following corollaries hold.
Corollary 2.4 (Gradient bias with second-order-smoothness).
Suppose $F:\mathbb{R}^{d}\to\mathbb{R}$ has partial derivates that are $L_{2}$-smooth w.r.t. $\|\cdot\|_{1}$. Let $x^{1},...,x^{T}\in\Delta_{d}$ be deterministic vectors and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for all $t\in[T]$. Then $\left\|\mathbb{E}\left[\nabla F\left(\bar{a}^{T}\right)-\nabla F\left(\bar{x}^{T}\right)\right]\right\|_{\infty}\leq\frac{2L_{2}}{T}$.
Proof 2.5.
This follows directly from applying Corollary 2.3 to each partial derivate: for all $j\in[d]$, we have $\left|\mathbb{E}\left[\nabla_{j}F\left(\bar{a}^{T}\right)-\nabla_{j}F\left(\bar{x}^{T}\right)\right]\right|\leq\frac{2L_{2}}{T}$.
Corollary 2.6 (Gradient bias with first-order-smoothness).
Suppose $F:\mathbb{R}^{d}\to\mathbb{R}$ is $L_{1}$-smooth w.r.t. $\|\cdot\|_{1}$.Let $x^{1},...,x^{T}$ be a collection of deterministic vectors in $\Delta_{d}$ and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for all $t\in[T]$. Then $\left\|\mathbb{E}\left[\nabla F\left(\bar{a}^{T}\right)-\nabla F\left(\bar{x}^{T}\right)\right]\right\|_{\infty}\leq\frac{4L_{1}}{\sqrt{T}}$.
Proof 2.7.
Denote $\bar{x}=\frac{1}{T}\sum_{t\in[T]}x^{t},\bar{a}=\frac{1}{T}\sum_{t\in[T]}a^{t}$. By $L_{1}$-smoothness of $F$, we have for all $j\in[d],r\in\mathbb{R}^{+}$: $F(\bar{x}+re_{j})\leq F(\bar{x})+\langle\nabla F(\bar{x}),re_{j}\rangle+\frac{L_{1}r^{2}}{2}\|e_{j}\|_{1}^{2}$. Re-arranging this expression gives $\langle\nabla F(\bar{x}),e_{j}\rangle\leq\frac{F(\bar{x})-F(\bar{x}+re_{j})}{r}+\frac{L_{1}r}{2}$. By $L_{1}$-smoothness of F-F, we can obtain $\langle\nabla F(\bar{a}),e_{j}\rangle\leq\frac{F(\bar{a}+re_{j})-F(\bar{a})}{r}+\frac{L_{1}r}{2}$. Adding these inequalities and taking expectations we get
$\mathbb{E}[\langle\nabla F(\bar{a})-\nabla F(\bar{x}),e_{j}\rangle]\leq\frac{\mathbb{E}\left[F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})\right]}{r}+L_{1}r.$
By applying Corollary 2.3 to $F(\cdot)$ and $F(\cdot+re_{j})$ we obtain
$\mathbb{E}\left[F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})\right]\leq 4L_{1}/T.$
Hence, for all $j\in[d]$, $\mathbb{E}[\nabla_{j}F(\bar{a})-\nabla_{j}F(\bar{x})]=\mathbb{E}[\langle\nabla F(\bar{a})-\nabla F(\bar{x}),e_{j}\rangle]\leq\frac{4L_{1}}{rT}+L_{1}r.$ Analogously we can prove the same bound for $\mathbb{E}[\langle\nabla F(\bar{x})-\nabla F(\bar{a}),e_{j}\rangle]$. Finally, choosing $r=2/\sqrt{T}$ gives the result.
We will also need to bound the second-moment of the error incurred by evaluating gradients at sampled vertices. Towards this, we first observe that the in-expectation guarantees obtained in Lemma 2.1 can be turned into high-probability guarantees by controlling the increments of the martingale difference sequence $(\alpha_{t})_{t\in[T]}$ with the bounded differences inequality.
Lemma 2.8 (Function value approximation with high probability).
Let $(\mathbf{E},\|\cdot\|)$ be a normed space and $F$ be an $L_{0}$-Lipschitz, $L_{1}$-smooth function w.r.t. $\|\cdot\|$. Let $x^{1},...,x^{T}$ be a collection of deterministic vectors, $(\lambda_{1},\ldots,\lambda_{T})\in\Delta_{T}$ and $a^{1},...,a^{T}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$. Suppose $\|x^{t}-a^{t}\|\leq D$ a.s. for all $t\in[T]$. Define $\bar{x}^{T}=\sum_{t=1}^{T}\lambda_{t}x^{t}$ and $\bar{a}^{T}=\sum_{t=1}^{T}\lambda_{t}a^{t}$. Then, for any $\beta>0$
$\mathbb{P}\left[F\left(\bar{a}^{T}\right)-F\left(\bar{x}^{T}\right)\geq\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}+\beta\sqrt{2}L_{0}D\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}\,\right]\leq\exp(-\beta^{2}).$
Proof 2.9.
Let $\alpha_{t}=\lambda_{t}\left\langle\nabla F\left(\sum_{k=t}^{T}\lambda_{k}x^{k}+\sum_{k=1}^{t-1}\lambda_{k}a^{k}\right),a^{t}-x^{t}\right\rangle$ as in the proof of Lemma 2.1. From that proof we have that $F(\bar{a}^{T})-F\left(\bar{x}^{T}\right)\leq\sum_{t=1}^{T}\alpha_{t}+\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}$. Hence
(4) $\mathbb{P}[F(\bar{a}^{T})-F\left(\bar{x}^{T}\right)\geq\alpha]\leq\mathbb{P}\left[\sum_{t=1}^{T}\alpha_{t}+\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\geq\alpha\right]=\mathbb{P}\left[\sum_{t\in[T]}\alpha_{t}\geq\alpha^{\prime}\right],$
where $\alpha=\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}+\alpha^{\prime}$. Next, we bound the probability on the right-hand side. Note that $(\alpha_{t})_{t\in[T]}$ is a martingale difference sequence and $|\alpha_{i}|\leq L_{0}D\lambda_{i}$, so the bounded differences inequality gives $\mathbb{P}\left[\sum_{t=1}^{T}\alpha_{t}\geq\beta\sqrt{2}L_{0}D\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}\right]\leq\exp(-\beta^{2})$, for any $\beta>0$. Making $\alpha^{\prime}=\beta\sqrt{2}L_{0}D\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}$ and replacing back into Equation 4, we conclude the claimed result.
We now use Lemma 2.8 to give bounds on the expected second-moment error, with respect to $\|\cdot\|_{\infty}$.
Lemma 2.10 (Second moment of maximum expected error).
Let $(\mathbf{E},\|\cdot\|)$ be a normed space and $F_{1},...,F_{M}$ a collection of $L_{0}$-Lipschitz and $L_{1}$-smooth functions w.r.t $\|\cdot\|$. Let $x^{1},\ldots,x^{T}\in\mathbb{R}^{d}$, $(\lambda_{1},\ldots,\lambda_{T})\in\Delta_{T}$ and $a^{1},...,a^{T}$ be a collection of random vectors such that $\mathbb{E}[a^{t}|a^{t-1},\ldots,a^{1}]=x^{t}$ and $\|a^{t}-x^{t}\|\leq D$ a.s. for all $t\in[T]$. Let $\bar{a}^{T}=\sum_{t\in[T]}\lambda_{t}a^{t}$ and $\bar{x}^{T}=\sum_{t\in[T]}\lambda_{t}x^{t}$. Then
$\mathbb{E}\left[\max_{j\in[M]}|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|^{2}\right]\leq\frac{L_{1}^{2}D^{4}}{2}\left(\sum_{t=1}^{T}\lambda_{t}^{2}\right)^{2}+2L_{0}^{2}D^{2}(4+\log(M))\sum_{t=1}^{T}\lambda_{t}^{2}.$
Proof 2.11.
Define the constants $A=\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}$ and $C=\sqrt{2}L_{0}D\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}$. We will introduce a quantity, $B$, in the following computation, and later optimize over the choice of this $B$. We proceed as follows:
$\mathbb{E}\bigg[\max_{j\in[M]}|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|^{2}$ $\bigg]=\int_{0}^{\infty}\mathbb{P}\left[\max_{j\in[M]}|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|\geq\beta\right]\beta d\beta$
$\hskip-65.44142pt\leq\int_{0}^{A+B}\beta d\beta+\sum_{j\in[M]}\int_{A+B}^{\infty}\mathbb{P}\left[|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|\geq\beta\right]\beta d\beta$
$\hskip-65.44142pt=\frac{(A+B)^{2}}{2}+\sum_{j\in[M]}\int_{B/C}^{\infty}\mathbb{P}\left[|F_{j}(\bar{x})-F_{j}(\bar{a}^{T})|\geq A+C\beta\right](A+C\beta)Cd\beta$
$\hskip-65.44142pt\leq\frac{(A+B)^{2}}{2}+2MC\left[C\int_{B/C}^{\infty}\exp\left(-\beta^{2}\right)\beta d\beta+A\int_{B/C}^{\infty}\exp\left(-\beta^{2}\right)d\beta\right]$
$\hskip-65.44142pt=\frac{(A+B)^{2}}{2}+2MC\left[C\exp\left(-(B/C)^{2}\right)+A\sqrt{\pi}\mathbb{P}[X\geq B/C]\right],$
where the first inequality comes from the union bound, the second from Lemma 2.8, and in the last line, we use density of a normal random variable, $X\sim{\cal N}(0,1/2)$. From Mill’s inequality, we have $\mathbb{P}[X\geq B/C]\leq\frac{C}{B\sqrt{\pi}}\exp\left(-(B/C)^{2})\right)$. Hence,
$\mathbb{E}\left[\max_{j\in[M]}|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|^{2}\right]\leq\frac{(A+B)^{2}}{2}+2MC(C+A)\exp\left(-(B/C)^{2}\right).$
Finally, setting $B=C\sqrt{\log(M)}$ if $M>1$ and $B=C$ if $M=1$ gives
$\mathbb{E}\left[\max_{j\in[M]}|F_{j}(\bar{a}^{T})-F_{j}(\bar{x}^{T})|^{2}\right]\leq 2A^{2}+(4+\log(M))C^{2}.$
We conclude this section by presenting two corollaries of Lemma 2.10.
Corollary 2.12 (Gradient approximation error with second-order-smoothness).
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be a $L_{1}$-smooth function w.r.t. $\|\cdot\|_{1}$ with $L_{2}$-smooth partial derivatives. Let $x^{1},...,x^{T}\in\Delta_{d}$ be a collection of deterministic vectors and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for all $t\in[T]$. Then
$\mathbb{E}\left\|\nabla F\left(\bar{a}^{T}\right)-\nabla F\left(\bar{x}^{T}\right)\right\|^{2}_{\infty}\leq\frac{8L_{2}^{2}}{T^{2}}+\frac{8L_{1}^{2}(4+\log(d))}{T}.$
Proof 2.13.
Apply Lemma 2.10 with $F_{j}=\nabla_{j}F$, $j\in[d]$.
Corollary 2.14 (Gradient approximation error with first-order-smoothness).
$\mathbb{E}\left\|\nabla F\left(\bar{a}^{T}\right)-\nabla F\left(\bar{x}^{T}\right)\right\|^{2}_{\infty}\leq\frac{8\sqrt{2}L_{1}^{2}}{T^{3/2}\sqrt{4+\log(d)}}+\frac{8\sqrt{2}(L_{0}^{2}+L_{1}^{2})\sqrt{(4+\log(d))}}{\sqrt{T}}.$
Proof 2.15.
Similarly to the proof of Corollary 2.6, using $L_{1}$-smoothness of $F$,
$\langle\nabla F(\bar{a})-\nabla F(\bar{x}),e_{j}\rangle$ $\leq\frac{F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})}{r}+L_{1}r$
$\leq\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})|}{r}+L_{1}r.$
Since F-F is also $L_{1}$-smooth, then
$\langle\nabla F(\bar{x})-\nabla F(\bar{a}),e_{j}\rangle\leq\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})|}{r}+L_{1}r.$
We conclude that
$\hskip-28.45274pt\leq 2\max_{j\in[d]}\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})|^{2}}{r^{2}}+L_{1}^{2}r^{2}.$
By Lemma 2.10 with $F_{j}(\cdot)=F(\cdot+re_{j})-F(\cdot)$ and $M=d$, we have
$\mathbb{E}\left[\max_{j\in[d]}\frac{|F(\bar{a}+re_{j})-F(\bar{x}+re_{j})+F(\bar{x})-F(\bar{a})|^{2}}{r^{2}}\right]\leq\frac{32L_{1}^{2}}{T^{2}r^{2}}+\frac{32L_{0}^{2}(4+\log(d))}{Tr^{2}}.$
Putting everything together and taking expectations
$\mathbb{E}[\left\|\nabla F\left(\bar{a}\right)-\nabla F\left(\bar{x}\right)\right\|^{2}_{\infty}]\leq\frac{64L_{1}^{2}}{T^{2}r^{2}}+\frac{64L_{0}^{2}(4+\log(d))}{Tr^{2}}+2L_{1}^{2}r^{2}.$
Choosing $r^{2}=\sqrt{\frac{32(4+\log(d))}{T}}$ gives the bound.
3 Differentially Private Stochastic Saddle Points
In this section we state and prove our main results regarding the DP-SSP problem in $\ell_{1}$ setting. Given a batch $B$ of iid samples from $\mathcal{D}$, we will denote $\nabla F_{\mathcal{D}}(x,y;B):=\frac{1}{|B|}\sum_{z\in B}\nabla f(x,y;z)$, and similarly for partial derivatives.
3.1 Private Stochastic Mirror Descent with Sampling from Vertices
Our first algorithm is a (entropic) stochastic mirror descent (SMD) method with sampling from vertices; see Algorithm 1. The following theorem holds.
Algorithm 1 Private stochastic mirror descent with vertex sampling
0: Dataset $S$, Number of steps $T$, Step-size $\tau$, Number of iterate samples $K$